# SAM Decoding: Speculative Decoding via Suffix Automaton

### v1

@@ 5 Related Work
Speculative Decoding. Speculative decoding is an approach that can significantly speed up large language models (LLMs) without compromising the quality of their outputs. This is achieved by increasing the number of tokens produced in each inference step while simultaneously decreasing the total number of inference steps required. The majority of speculative decoding techniques rely on smaller neural networks to create initial drafts during the inference process. These techniques are referred to as model-based speculative decoding methods. Early implementations of model-based speculative decoding, such as those described in Speculative Decoding (Leviathan et al., 2023), primarily focused on generating draft sequences using pre-existing, smaller-scale LLMs. Subsequently, advancements like Medusa (Cai et al., 2024) and SpecInfer (Miao et al., 2024) introduced tree-based speculative methods and initiated the development of dedicated draft models tailored for speculative decoding. For instance, Medusa developed a set of specialized decoding heads designed to serve as draft models. More recently, EAGLE (Li et al., 2024a, b) highlighted the significance of incorporating the hidden state of the model’s output into the speculative decoding process. In line with this insight, EAGLE trained a draft model that integrates the hidden state information from the model’s output, leading to more efficient acceleration across a range of tasks.
@@
In contrast to model-based methods, certain approaches focus on generating drafts by leveraging the model’s recent prediction outcomes, such as the output from the decoding head. Examples of such methods include Lookahead Decoding (Fu et al., 2024) and Token Recycling (Luo et al., 2024). Lookahead Decoding keeps track of the $n$-grams of the generated text, whereas Token Recycling maintains the top $k$ probable next tokens for each tokens.
@@
Moreover, some strategies generate drafts directly through the retrieval of previously generated texts or text corpora, utilizing $n$-gram matching. Notable among these are PLD (Saxena, 2023) and REST (He et al., 2024). Our proposed SAM-Decoding method also employs retrieval for draft generation. However, it stands out due to its superior speed and accuracy, and it can be seamlessly integrated with other speculative decoding techniques.
@@
Additionally, beyond the aforementioned methods, research also conducted on speculative decoding that relies either on the model itself (Kou et al., 2024) or on sub-models within the larger architecture (Elhoushi et al., 2024).
@@
Efficient LLM Architecture. There is also work to improve the model’s inference speed from the perspective of model structure. This part of the work includes model distillation, quantization and pruning. Model distillation (Sreenivas et al., 2024; Muralidharan et al., 2024) distills the knowledge of a large model into a small model, thereby speeding up inference while maintaining the model’s performance. Quantization (Frantar et al., 2022; Xiao et al., 2023; Lin et al., 2024; Liu et al., 2024; Ashkboos et al., 2024b) reduces the number of bits required to store parameters and reduces the data transmission time from HBM to on-chip memory during inference, thereby achieving effective inference acceleration. Pruning (Frantar & Alistarh, 2023; Ashkboos et al., 2024a; Men et al., 2024; Chen et al., 2024; Hu et al., 2024; Sun et al., 2024; Zhang et al., 2024) is used to remove unimportant parameters in the model. For structured pruning, it can be combined with model distillation to train efficient small models, while semi-structured pruning can reduce the model’s memory access and computing overhead and improve the inference speed by combining special hardware, i.e., sparse tensor core.
@@ A.1 Construction Process of Suffix Automaton
A.2 Time Complexity of Suffix Automaton
Consider a suffix automaton $S$ with an initial state $s_{0}$, which corresponds to the root node of the automaton (representing the empty string). Suppose that state $s_{0}$ undergoes transitions through a sequence of $L$ tokens $x=(x_{1},x_{2},\ldots,x_{L})$:
@@ A.2 Time Complexity of Suffix Automaton
First, let us define the matching length associated with state $s_{i}$ as $l_{i}$. Given that each state transition can increase the matching length by at most 1, it follows that $0\leq l_{i}\leq i$. Next, we introduce the concept of energy $\phi$ for each state $s_{i}$, defined as $\phi(s_{i})=l_{i}$. Let $c_{i}$ represent the time cost of the $i$-th state transition. We then define the amortized cost $\hat{c_{i}}$ as:
@@
Next, we analyze the upper bound of $\hat{c_{i}}$. Each state transition involves moving through the link edge zero or more times, followed by a move through the next edge. Transitioning via the link edge incurs a cost of 1 but decreases the potential by at least 1. Conversely, transitioning via the next edge also incurs a cost of 1 but increases the potential by 1. Consequently, the amortized cost $\hat{c_{i}}$ is bounded above by 2, leading to:

### v2

@@ 5 Related Work
Speculative Decoding. Speculative decoding is an approach that can significantly speed up large language models (LLMs) without compromising the quality of their outputs. This is achieved by increasing the number of tokens produced in each inference step while simultaneously decreasing the total number of inference steps required. The majority of speculative decoding techniques rely on smaller neural networks to create initial drafts during the inference process. These techniques are referred to as model-based speculative decoding methods. Early implementations of model-based speculative decoding, such as those described in Speculative Decoding Leviathan et al. (2023), primarily focused on generating draft sequences using pre-existing, smaller-scale LLMs. Subsequently, advancements like Medusa Cai et al. (2024) and SpecInfer Miao et al. (2024) introduced tree-based speculative methods and initiated the development of dedicated draft models tailored for speculative decoding. For instance, Medusa developed a set of specialized decoding heads designed to serve as draft models. More recently, EAGLE Li et al. (2024b, a) highlighted the significance of incorporating the hidden state of the model’s output into the speculative decoding process. In line with this insight, EAGLE trained a draft model that integrates the hidden state information from the model’s output, leading to more efficient acceleration across a range of tasks.
@@
In contrast to model-based methods, certain approaches focus on generating drafts by leveraging the model’s recent prediction outcomes, such as the output from the decoding head. Examples of such methods include Lookahead Decoding Fu et al. (2024) and Token Recycling Luo et al. (2024). Lookahead Decoding keeps track of the $n$-grams of the generated text, whereas Token Recycling maintains the top $k$ probable next tokens for each tokens.
@@
Moreover, some strategies generate drafts directly through the retrieval of previously generated texts or text corpora, utilizing $n$-gram matching. Notable among these are PLD Saxena (2023) and REST He et al. (2024). Our proposed SAM-Decoding method also employs retrieval for draft generation. However, it stands out due to its superior speed and accuracy, and it can be seamlessly integrated with other speculative decoding techniques.
@@
Additionally, beyond the aforementioned methods, research also conducted on speculative decoding that relies either on the model itself Kou et al. (2024) or on sub-models within the larger architecture Elhoushi et al. (2024).
@@
Efficient LLM Architecture. There is also work to improve the model’s inference speed from the perspective of model structure. This part of the work includes model distillation, quantization and pruning. Model distillation Sreenivas et al. (2024); Muralidharan et al. (2024) distills the knowledge of a large model into a small model, thereby speeding up inference while maintaining the model’s performance. Quantization Frantar et al. (2022); Xiao et al. (2023); Lin et al. (2024); Liu et al. (2024); Ashkboos et al. (2024b) reduces the number of bits required to store parameters and reduces the data transmission time from HBM to on-chip memory during inference, thereby achieving effective inference acceleration. Pruning Frantar and Alistarh (2023); Ashkboos et al. (2024a); Men et al. (2024); Chen et al. (2024); Hu et al. (2024); Sun et al. (2024); Zhang et al. (2024) is used to remove unimportant parameters in the model. For structured pruning, it can be combined with model distillation to train efficient small models, while semi-structured pruning can reduce the model’s memory access and computing overhead and improve the inference speed by combining special hardware, i.e., sparse tensor core.
@@ A.1 Construction Process of Suffix Automaton
Algorithm 3 introduces the construction (Build-SAM) and expansion process (Expand) of Suffix Automaton, where INIT_SAM function will create a suffix automaton that only contains the root node. And for the root node, the link attribute value is $-1$, the next attribute value is empty, the length attribute value is 00, and the min_endpos attribute value is 00.
@@
A.2 Time Complexity of State Transfer
In this section, we introduce the time complexity of state transfer of suffix automaton. Consider a suffix automaton $S$ with initial state $s_{0}$, which corresponds to the root node of the automaton (representing the empty string). Suppose that state $s_{0}$ undergoes transitions through a sequence of $L$ tokens $x=(x_{1},x_{2},\ldots,x_{L})$:
@@ A.2 Time Complexity of State Transfer
First, let us define the matching length associated with state $s_{i}$ as $l_{i}$. Given that each state transition can increase the length of the match by at most 1, it follows that $0\leq l_{i}\leq i$. Next, we introduce the concept of energy $\phi$ for each state $s_{i}$, defined as $\phi(s_{i})=l_{i}$. Let $c_{i}$ represent the time cost of the transition of the $i$-th state. We then define the amortized cost $\hat{c_{i}}$ as:
@@
Next, we analyze the upper bound of $\hat{c_{i}}$. Each state transition involves moving through the link edge zero or more times, followed by a move through the next edge. Transitioning through the link edge incurs a cost of 1 but decreases the potential by at least 1. Conversely, transitioning through the next edge also incurs a cost of 1 but increases the potential by 1. Consequently, the amortized cost $\hat{c_{i}}$ is bounded above by 2, leading to:
@@
Llama3-8B-Inst PLD 1.30$\times$ 1.12$\times$ 1.41$\times$ 1.03$\times$ 1.30$\times$ 1.53$\times$ $1.39$ 44.26 1.28$\times$
Token Recycling 1.92$\times$ 1.88$\times$ 1.85$\times$ 1.75$\times$ 2.24$\times$ 1.82$\times$ 2.76 66.42 1.91$\times$
SAM-Decoding[T] 2.09$\times$ 1.93$\times$ 2.04$\times$ 1.82$\times$ 2.32$\times$ 2.12$\times$ 2.63 85.73 2.05$\times$
EAGLE2 2.08$\times$ 1.95$\times$ 1.85$\times$ 1.80$\times$ 2.31$\times$ 1.87$\times$ 3.90 68.69 1.98$\times$
SAM-Decoding[E2] 2.23$\times$ 1.92$\times$ 1.93$\times$ 1.76$\times$ 2.23$\times$ 2.06$\times$ 3.91 70.40 2.03$\times$
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Llama3-8B-Inst PLD 1.30 42.39 1.18×1.18$\times$1.18 × 1.23 39.19 1.06×1.06$\times$1.06 × 1.50 45.15 1.34×1.34$\times$1.34 ×
Token Recycling 2.93 71.49 1.99×1.$9$ 2.89 74.05 2.02×2.02$\times$2.02 × 2.84 62.77 1.87×1.87$\times$1.87 ×
SAM-Decoding[T] 2.77 78.04 2.16×2.16$\times$2.16 × 2.65 76.76 2.09×2.09$\times$2.09 × 2.70 66.76 1.99×1.$9$
EAGLE2 4.74 85.58 2.37×2.37$\times$2.37 × 4.77 95.02 2.58×2.58$\times$2.58 × 3.97 63.30 1.86×1.86$\times$1.86 ×
SAM-Decoding[E2] 4.76 91.50 2.54×2.54$\times$2.54 × 4.34 89.15 2.43×2.43$\times$2.43 × 3.93 67.94 2.02×2.02$\times$2.02 ×
Vicuna-13B PLD 1.61$\times$ 1.10$\times$ 2.36$\times$ 1.11$\times$ 1.69$\times$ 1.80$\times$ 1.66 33.89 1.59$\times$
Token Recycling 2.03$\times$ 1.84$\times$ 2.07$\times$ 1.83$\times$ 2.42$\times$ 1.84$\times$ 2.81 42.74 2.01$\times$
SAM-Decoding[T] 2.36$\times$ 1.80$\times$ 2.63$\times$ 1.83$\times$ 2.49$\times$ 2.22$\times$ 2.91 47.27 2.22$\times$
EAGLE2 3.10$\times$ 2.15$\times$ 2.58$\times$ 2.38$\times$ 3.19$\times$ 2.33$\times$ 4.42 56.06 2.63$\times$
SAM-Decoding[E2] 3.27$\times$ 2.12$\times$ 2.89$\times$ 2.34$\times$ 3.12$\times$ 2.54$\times$ 4.51 57.88 2.72$\times$
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-13B PLD 1.54 32.06 1.44×1.$4$ 1.40 29.43 1.30×1.30$\times$1.30 × 1.90 43.38 2.15×2.15$\times$2.15 ×
Token Recycling 2.79 46.03 2.07×2.07$\times$2.07 × 2.85 48.89 2.16×2.16$\times$2.16 × 2.90 40.97 2.03×2.03$\times$2.03 ×
SAM-Decoding[T] 2.79 50.87 2.28×2.28$\times$2.28 × 2.83 52.14 2.30×2.30$\times$2.30 × 2.99 48.33 2.40×2.40$\times$2.40 ×
EAGLE2 5.15 77.85 3.49×3.49$\times$3.49 × 5.26 82.13 3.63×3.63$\times$3.63 × 4.24 52.28 2.59×2.59$\times$2.59 ×
SAM-Decoding[E2] 4.97 73.78 3.31×3.31$\times$3.31 × 4.80 73.94 3.27×3.27$\times$3.27 × 4.41 56.17 2.78×2.78$\times$2.78 ×
Vicuna-33B PLD 1.50$\times$ 1.07$\times$ 2.06$\times$ 1.09$\times$ 1.59$\times$ 1.51$\times$ 1.65 13.33 1.46$\times$
Token Recycling 2.10$\times$ 1.84$\times$ 2.19$\times$ 1.88$\times$ 2.42$\times$ 1.92$\times$ 2.70 18.8 2.06$\times$
SAM-Decoding[T] 2.31$\times$ 1.79$\times$ 2.53$\times$ 1.90$\times$ 2.48$\times$ 2.06$\times$ 2.68 19.87 2.18$\times$
EAGLE2 3.29$\times$ 2.31$\times$ 2.73$\times$ 2.51$\times$ 3.65$\times$ 2.46$\times$ 4.06 25.86 2.83$\times$
SAM-Decoding[E2] 3.36$\times$ 2.24$\times$ 2.88$\times$ 2.41$\times$ 3.43$\times$ 2.50$\times$ 4.08 25.60 2.81$\times$
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-33B PLD 1.58 14.18 1.51×1.51$\times$1.51 × 1.34 11.82 1.30×1.30$\times$1.30 × - - -
Token Recycling 2.64 19.64 2.09×2.09$\times$2.09 × 2.76 20.92 2.19×2.19$\times$2.19 × - - -
SAM-Decoding[T] 2.73 22.44 2.39×2.39$\times$2.39 × 2.65 21.80 2.28×2.28$\times$2.28 × 2.60 19.74 2.26×2.26$\times$2.26 ×
EAGLE2 3.53 28.18 3.00×3.$0$ 5.04 37.35 3.91×3.91$\times$3.91 × 3.84 24.28 2.78×2.78$\times$2.78 ×
SAM-Decoding[E2] 3.61 29.49 3.14×3.14$\times$3.14 × 4.70 34.21 3.58×3.58$\times$3.58 × 3.79 24.45 2.80×2.80$\times$2.80 ×
Vicuna-7B PLD 1.61$\times$ 1.56$\times$ 0.99$\times$ 2.47$\times$ 1.11$\times$ 1.51$\times$ 1.75 78.26 1.56$\times$
Token Recycling 2.08$\times$ 1.76$\times$ 1.97$\times$ 1.85$\times$ 2.35$\times$ 1.76$\times$ 2.82 98.39 1.96$\times$
SAM-Decoding[T] 2.62$\times$ 1.82$\times$ 2.92$\times$ 2.09$\times$ 2.60$\times$ 2.21$\times$ 3.02 119.21 2.38$\times$
EAGLE2 2.66$\times$ 1.76$\times$ 2.18$\times$ 2.03$\times$ 2.63$\times$ 1.97$\times$ 4.34 110.56 2.21$\times$
SAM-Decoding[E2] 3.19$\times$ 1.97$\times$ 2.86$\times$ 2.28$\times$ 2.84$\times$ 2.32$\times$ 4.52 129.36 2.58$\times$
Table 10: Speedup of SAM-Decoding on A100 GPU compared to the baselines on Spec-Bench.
Appendix B Additional Experiment Results
In this section, we present the results of the experiment on Llama3-8B-instruct.
Tables 4 and 5 present the speedup ratios of SAM-Decoding compared to baseline methods across the Spec-Bench, HumanEval, MBPP, and HAGRID datasets, utilizing the Llama3-8B-instruct model. In particular, SAM-Decoding , when paired with Token Recycling (SAM-Decoding[T]), brings speedups on all tasks and achieves the highest average speedup. Specifically, SAM-Decoding enhances the speedup ratio of Token Recycling from 1.92$\times$, 1.85$\times$, and 1.82$\times$ to 2.09$\times$, 2.04$\times$, and 2.12$\times$ for multi-turn conversation, summarization, and retrieval-augmented generation tasks, respectively. This improvement raises the overall speedup ratio of token recycling in the Spec-Bench dataset from 1.91$\times$ to 2.05$\times$. On the HumanEval and MBPP datasets, SAM-Decoding increases the speedup ratio of Token Recycling from 1.80$\times$ and 2.04$\times$ to 2.27$\times$ and 2.37$\times$, respectively. For the HAGRID dataset, the integration of SAM-Decoding with Token Recycling increases the speedup ratio from 1.87$\times$ to 1.99$\times$. Furthermore, beyond enhancing Token Recycling, SAM-Decoding also amplifies the performance gains of EAGLE2 in multi-turn conversation, summarization, retrieval-augmented generation, and context Q&A tasks. The speedup ratios were increased from 2.08$\times$, 1.85$\times$, 1.87$\times$, and 1.86$\times$ to 2.23$\times$, 1.93$\times$, 2.06$\times$, and 2.02$\times$ respectively.
Appendix C Additional Ablation Experiments
Figure 4: the percentage of inference time of different modules in SAM-Decoding.
Figure 5: the percentage of usage times and mean accept tokens of different draft modules.
In this section, we present additional ablation experiments, including the percentage of inference time of different modules in the decoding process of SAM-Decoding, and the percentage of drafts provided by different draft modules in SAM-Decoding.
The inference process of SAM-Decoding is divided into five stages: prefill, draft generation, decoding, verification, and updating. During the prefill stage, the model processes the input prompt to establish an initial state. In the draft generation stage, a draft is produced based on this initial state. The decoding stage involves the model further processing this draft. Next comes verification, where the correct parts of the draft are evaluated based on the information processed during the decoding stage. Finally, the update phase modifies the state of the model based on the valid parts of the draft. Figure 4 illustrates the proportion of time each stage consumes within the SAM-Decoding[T] process base on Spec-Bench. As shown, the decoding stage takes up the largest portion of time, accounting for 65.4% of the entire process. This is followed by the verification stage, which occupies 23.4% of the total time. The updating stage requires 6.3% of the time, whereas the draft generation stage contributes only 0.6% to the overall duration. Additionally, the prefill stage comprises 4.2% of the total processing time.
Figure 5 shows the usage times of different draft modules of SAM-Decoding[T] on Spec-Bench and the corresponding average draft accept length. It can be seen that in 85.96% of the cases, due to insufficient matching length, we generate drafts based on the auxiliary method, corresponding to an average accept length of 2.51, while in the remaining 11.59% and 2.45% of the cases, the dynamic suffix automaton and static suffix automaton are used to generate drafts, corresponding to average accept lengths of 6.57 and 3.39, respectively.